show that the straight line (a+2b)x+(a-3b)y+b-a=0 always passes through a fixed point .find coordinates of that fixed point. i need the answer urgently plz😖
 
 
 
 
 
 
 
 
 
 

Grade:11

1 Answers

Arun
6 years ago
Dear student
(x+y-1)a + (2x-3y+1) b = 0
Hence
x +y-1 = 0
x +y = 1  and  2x -3y = -1
On solving
5x = 2
x = 2/5
and 
y = 3/5
Regards
Arun (askIITians forum expert)